-3y+6x= -18 what are the intercepts
Answer by algebrahouse.com(1659)

To find the x-intercept, put 0 in for y
-3y + 6x = -18
-3(0) + 6x = -18 {substituted 0 in for y}
6x = -18 {multiplied -3 by 0}
x = -3 {divided both sides by 6}
x-intercept is -3
coordinates are (-3,0)

To find they-intercept, put 0 in for x
-3y + 6x = -18
-3y + 6(0) = -18 {substituted 0 in for x}
-3y = -18 {multiplied 6 by 0}
y = 6 {divided both sides by -3}
y-intercept is 6
coordinates are (0,6)

For more help from me, visit: www.algebrahouse.com